The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by checkley71(8403)

2L+2W=P
L=2W-5
2(2W-5)+2W=230
4W-10+2W=230
6W=230+10
6W=240
W=240/6
W=40 SOLUTION
L=2*40-5
L=80-5
L=75 SOLUTION
PROOF
2*75+2*40=230
150+80=230
230=230